OpenStudy (anonymous):

what are parametric equations direction numbers, symmetric equations, and normal vectors ? what are parametric equations direction numbers, symmetric equations, and normal vectors ? @MIT 18.02 Multiva…

OpenStudy (anonymous):

and what are they used to find

OpenStudy (anonymous):

PARAMETRIC EQUATIONS are used to find points on a region they are also used to find velocity and acceleration they are in the form x= x(t), y=y(t), z=z(t) in order to convert two parametric equations to one, 1. if you can, elimate t, if not good luck {by solving for t} 2.sub expression into other thhhhhhhheeeeere yaaaaaaaaaa goooooo

OpenStudy (anonymous):

positive direction of a vector is from p1-p2 negative direction of a vector is from p2-p1 Direction Numbers are any 3#s proportional to direction cosiness of a line seg(vector) Direction angles they are the angles that P1P2 (vector) make with pos x, y,z coord axis[alpha, beta, weird y resp.] Direction cosines - cosines of direction angles using two points P1P2 ex. cos alpha= xsub2-xsub1 , cos beta =ysub2 -ysub1, cos weird y=zsub2-zsub1 {separately each is divided by d [distance bewteen P1P2=length-magnitude note that cos ^2alpha +cos^2beta+cos^2 weird y=1 , lambda, curs u, curs,v resp.

OpenStudy (anonymous):

symmetric equations a line needs a pt and a direction *note should be at top you can get xyz intercepts from equation of plane by setting the other two variables =o aand you can also get perpendicular -orthogonal-normal direction by taking the cooeffiec of x y and z from equation of plane *note parallel vectors r usually given, perp orth norm. u have to find take pt and direction plug into r(t)[vector equation]=rsub 0(point) + t *v(direction from xyz coeff. of eqaution of plane) *note parametric equations are derived [find symmetricla equations, or you can find points on the line by plugging in any number] for symmet equations solve each paramtetric equation for t thus symmetrical equations= each t parameter = to each you can use symmetric equations to find where a line intersects a certain plane by setting the left out variable =0 and solving for the plane variables, giving you a point( ) where left out=0 to show that lines are skew given only parametric equations refer to parametric equation the corresponding t variables are vectors <> using these vectors the vectors will be parallel if scalar multiple sof eahc other. the vectors will intersect if you can take line 1 parametric equationx=line 2 parametric eqauationx, same with remainding parametric equations , then solve for s and t and plg into remaining l1=l2 param. equation. if you obtain a true statement they intersect. those lines are perpendicular orth normal if dot prodct =0 *check by subbing in any unit vector. function of two variables is skew-symmetric if f(y, x) = −f(x, y). The property implies f(x, x) = 0 (except in fields of characteristic two).